# MAGNET: an open-source library for mesh agglomeration by Graph Neural Networks

### v1

@@ 10 Data, SimType, model, mode, param);
Figure 15: Verification Test 1: PolyDG solution of problem (8) computed on a grid of 256 agglomerated elements (left), plot of the exact solution (center) and computed numerical error (right).
@@
Figure 16: Verification Test 1: Computed errors $||u_{h}-u_{ex}||_{L^{2}(\Omega)}$ and $||u_{h}-u_{ex}||_{dG}$ as a function of the mesh size $h$ by fixing the polynomial degree $\ell=3$. On the left, the test is performed with k-means by agglomerating different meshes so that the ratio of original elements to agglomerated elements is the same; on the right, the test is performed by agglomerating the same mesh with SAGE-Base using an increasing number of recursive bisections.
@@ 5.2 Verification Test 2: Heat Equation with Discontinuous Boundary Conditions
We solve (9) through lymph, by combining a PolyDG method with a Crank-Nicolson time integration scheme, [27]. We consider $\Omega=B_{0.6}(0.5,0)\cup B_{0.6}(-0.5,0)$, i.e., the combination of two balls that intersect. We take $\mu=0.1$, final time $T=1$ and discontinuous Dirichlet boundary conditions $u(\mathbf{x})=1$ for $\mathbf{x}_{1}\geq 0$, $u(\mathbf{x})=0$ for $\mathbf{x}_{1}<0$, with analogous discontinuous initial condition. We solve the problem on three meshes agglomerated by METIS, k-means, and SAGE-Base, respectively, starting from a very fine mesh of 35158 elements, using either k-way or number of refinements mode so that the resulting mesh has 256 elements. Additionally, we consider for comparison a polygonal mesh with the same number of elements generated using the software Polymesher [52], which tends to create very regular hexagonal cells. For each mesh, the problem is discretized using 3 different polynomial degrees $\ell=1,3,5$ and time step $\Delta t=10^{-3}$. The computed PolyDG solutions at the final time $T=1$ are reported in Figure 17.
@@
We observe that all four meshes produce very similar results when considering the same polynomial degree, although the agglomerated meshes have a significantly higher number of edges per element compared to the Polymesher one. The numerical solutions are overall coherent with what is expected from the literature [53]. For higher polynomial degrees, some numerical oscillations appear in proximity to the boundary conditions discontinuity; this might be due to agglomerated elements not perfectly conforming to the two different sections of the border. Indeed, if a single element shares an edge with both sections, there will be great numerical oscillations because a continuous polynomial would need to approximate a discontinuous function. Unfortunately, MAGNET has no way to guarantee this does not happen since the boundary is not taken into account when agglomerating.
@@ Acknowledgements.
We thank Mattia Corti for segmenting and providing the meshes of the brain. The brain MRI images were provided by OASIS-3: Longitudinal Multimodal Neuroimaging: Principal Investigators: T. Benzinger, D. Marcus, J. Morris; NIH P30 AG066444, P50 AG00561, P30 NS09857781, P01 AG026276, P01 AG003991, R01 AG043434, UL1 TR000448, R01 EB009352. This work received funding from the European Union (ERC SyG, NEMESIS, project number 101115663). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. PFA, MC and IM are members of INdAM-GNCS. The present research is part of the activities of “Dipartimento di Eccellenza 2023-2027”, MUR, Italy.
@@
The code and data used in this work are available can be accessed at https://github.com/lymphlib/magnet. Proper attribution should be given when reusing or distributing the materials.

### v2

@@ 10 Data, SimType, model, mode, param);
Figure 17: Verification Test 1.1: Computed errors $u_{h}-u_{ex}||_{L^{2}(\Omega)}$ and $u_{h}-u_{ex}||_{dG}$ as a function of the square root of degrees of freedom for problem (10) solved on an agglomerate triangular grid of the unit square of $2\cdot 1024^{2}$ elements with different agglomeration methods.
Figure 18: Verification Test 1.2: PolyDG solution of problem (10) computed on a grid of 256 agglomerated elements (left), plot of the exact solution (center) and computed numerical error (right).
Figure 19: Verification Test 1.2: Computed errors $u_{h}-u_{ex}||_{L^{2}(\Omega)}$ and $u_{h}-u_{ex}||_{dG}$ as a function of the mesh size $h$ by fixing the polynomial degree $\ell=3$. On the left, the test is performed with k-means by agglomerating different meshes so that the ratio of original elements to agglomerated elements is the same; on the right, the test is performed by agglomerating the same mesh with SAGE-Base using an increasing number of recursive bisections.
@@ 5.2 Verification Test 2: Heat Equation with Discontinuous Boundary Conditions
We solve (11) through lymph, by combining a PolyDG method with a Crank-Nicolson time integration scheme, [27]. We consider $\Omega=B_{0.6}(0.5,0)\cup B_{0.6}(-0.5,0)$, i.e., the combination of two balls that intersect. We take $\mu=0.1$, final time $T=1$ and discontinuous Dirichlet boundary conditions $u(\mathbf{x})=1$ for $\mathbf{x}_{1}\geq 0$, $u(\mathbf{x})=0$ for $\mathbf{x}_{1}<0$, with analogous discontinuous initial condition. We solve the problem on three meshes agglomerated by METIS, k-means, and SAGE-Base, respectively, starting from a very fine mesh of 35158 elements, using either k-way or number of refinements mode so that the resulting mesh has 256 elements. Additionally, we consider for comparison a polygonal mesh with the same number of elements generated using the software Polymesher [52], which tends to create very regular hexagonal cells. For each mesh, the problem is discretized using 3 different polynomial degrees $\ell=1,3,5$ and time step $\Delta t=10^{-3}$. The computed PolyDG solutions at the final time $T=1$ are reported in Figure 20. We observe that all four meshes produce very similar results when considering the same polynomial degree, although the agglomerated meshes have a significantly higher number of edges per element compared to the Polymesher one. The numerical solutions are overall coherent with what is expected from the literature [53]. For higher polynomial degrees, some numerical oscillations appear in proximity to the boundary conditions discontinuity; this might be due to agglomerated elements not perfectly conforming to the two different sections of the border. Unfortunately, our current architecture does not allow for taking into account information belonging to the frontier of a domain, and it is an active topic of research. Indeed, if a single element shares an edge with both sections, there will be great numerical oscillations because a continuous polynomial would need to approximate a discontinuous function. Unfortunately, MAGNET has no way to guarantee this does not happen since the boundary is not taken into account when agglomerating.
@@ 6 Conclusions
Declarations
Funding.
This work received funding from the European Union (ERC SyG, NEMESIS, project number 101115663). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ Acknowledgements.
We thank Mattia Corti for segmenting and providing the meshes of the brain. The brain MRI images were provided by OASIS-3: Longitudinal Multimodal Neuroimaging: Principal Investigators: T. Benzinger, D. Marcus, J. Morris; NIH P30 AG066444, P50 AG00561, P30 NS09857781, P01 AG026276, P01 AG003991, R01 AG043434, UL1 TR000448, R01 EB009352. Paola F. Antonietti, Matteo Caldana and Ilario Mazzieri are members of INdAM-GNCS. The present research is part of the activities of “Dipartimento di Eccellenza 2023-2027”, MUR, Italy.
@@
The code and data used in this work are available and can be accessed at https://github.com/lymphlib/magnet. Proper attribution should be given when reusing or distributing the materials.
Authors’ contribution.
P.F.A.: Conceptualization, Funding acquisition, Methodology, Project administration, Supervision, Writing - review and editing. M.C.: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Validation, Visualization, Writing - review and editing. I.M.: Conceptualization, Methodology, Project administration, Supervision, Writing - review and editing. A.R.F.: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Validation, Visualization, Writing - original draft.